# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v2

@@ 3.2 Infrared
The Palomar 200-inch (P200) near-infrared telescope equipped with the Wide Field Infrared Camera (WIRC, Wilson et al. 2003) observed AT2019aalc on 2023-07-05. The imaging was obtained in the J-, H,- and Ks-bands. The data was processed with the Swarp software (Bertin et al. 2002). We performed background subtraction using the photutils.background class and measured the flux using a circular aperture with a radius of pixels. The corresponding zero-point magnitudes for this aperture size were calculated using the automatic pipeline for each filter. The error bars are the computed Poisson and background error of the photometry222https://www.stsci.edu/hst/instrumentation/wfc3/data-analysis/photometric-calibration/ir-photometric-calibration and we also took into account the uncertainty of the derived zero-point magnitudes.
We obtained mid-infrared photometry from archival observations of the Wide-Field Survey Explorer (WISE, Wright et al. 2010), which was continuously monitoring the sky at (W1) and, due to the NEOWISE project (Mainzer et al. 2011), at (W2). It visits any point in the sky roughly every six months for about ten times. Using timewise (Necker & Mechbal 2024; Necker et al. 2024), we downloaded all available multi-epoch photometry data from the AllWISE data release and single-epoch photometry from the NEOWISE-R data releases and stacked the data per visit to obtain robust measurements. All selected single-exposure photometry data points are within arcsecond of the nucleus of the host. The resulting light curve spans almost years, starting in February until July 2024.
@@ 3.3 X-ray
The X-ray Telescope (XRT, Burrows et al. 2005) observations were obtained simultaneously with the UVOT observations. The observations were performed in photon counting mode. We performed the data reduction using the Space Science Data Center (SSDC) interactive archive in the following way. First, cleaned event-files were produced using the xrtpipeline and xrtexpomap tasks. The ximage tool allowed us to determine the source detection level and its position. In case of non-detections, upper limits for the count rate were derived and converted to flux using WebPIMMS333https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl and assuming the mean power-law index of the observations that resulted in detections. When our target source was detected above the level, we ran swxrtdas in order to extract the source and the background spectra. Source counts were retrieved from a circle with a radius of pixel ( arcseconds), while background counts were extracted using an annular region with pixel inner/outer radius. Finally, the grppha tool of the heasoft 6.32 software package was used to perform background subtraction using the source and background spectra. We binned each spectrum to have a minimum of counts per spectral bin.
We fit the individual X-ray spectra using the XSPEC fitting environment (Arnaud 1996b). First, we have determined the Galactic neutral hydrogen column density, , based on the HI4PI survey (HI4PI Collaboration et al. 2016). The spectrum was then fitted using Cash-statistic (C-stat, Cash 1976). To find the best-fitting model, we first fitted the stacked spectrum with a absorbed power-law (Tbabs*powerlaw in XSPEC) model as the last XMM data of AT2019aalc before our monitoring program was best-fit with a power-law model (Pasham 2023). When fitting the model, the galactic column density was fixed whilst the power-law index and the normalization were free to vary. The resultant fit quality is C-Stat for degrees of freedom. We than added a blackbody model component, to take into account a possible soft excess. The blackbody temperature and the normalization of the new component were also free to vary and we fit again which resulted in a better fit with C-Stat for degrees of freedom. We implemented the likelihood ratio test (lrt in XSPEC) to estimate the statistical significance of the inclusion of the blackbody component. After running simulations, the test resulted only in smaller likelihood differences than the observed one clearly implying that the difference is unlikely to be due to random fluctuations. Therefore, accounting for the soft X-ray excess results in a significantly better fit.
@@ 3.4.1 VLBI observations
We observed AT2019aalc with the European VLBI Network (EVN) and the enhanced Multi-Element Remotely Linked Interferometer Network (e-MERLIN, Garrington et al. 2004). These observations were carried out at GHz (L-band) on 2023 June 09 (project code: EV027, PI: Veres). The observation time was hr. The experiment was performed in phase-referencing mode with -min duty cycles (with spending min on the target during each cycle) which resulted in an on-source time of h including fringe-finder scans and slewing. The data were recorded in left and right circular polarizations with a data rate of Gbit s-1. The correlation was done at the EVN Data Processor (Keimpema et al. 2015) at the Joint Institute for VLBI European Research Infrastructure Consortium (Dwingeloo, The Netherlands) with an averaging time of s. The Compact Symmetric Object (CSO) 15210430 (a.k.a. 4C 04.51) at an angular distance of was observed as phase-reference calibrator.
The data were calibrated with the U.S. National Radio Astronomy Observatory Astronomical Image Processing System (AIPS, Greisen 1990) software package following the standard procedures including time flagging, amplitude calibration, correction for the dispersive ionospheric delays and global fringe-fitting of the calibrator sources. We imaged the phase-calibrator with the Difmap software (Shepherd et al. 1994), using the hybrid mapping technique (Högbom 1974). Using the clean component (CC) model we built up in Difmap, we repeated the fringe-fitting in AIPS for the phase-reference calibrator source to correct its structure. Finally, we interpolated the solutions of the last fringe-fitting to the data of our target source. We imaged the target source using the hybrid mapping technique in Difmap. At the end of the procedure, several steps of amplitude and phase self-calibration were performed. A simple elliptical Gaussian modelfit component (Pearson 1995) was fitted directly to the visibility data. Finally, to decrease the noise level, we performed clean iterations in the residual image with a small loop gain of .
@@ 3.4.2 ATCA monitoring
The Australia Telescope Compact Array (ATCA)444https://csiropedia.csiro.au/australia-telescope-compact-array/, located at Narrabri, New South Wales, Australia, was used to monitor AT2019aalc. The observations were conducted between 2023-06-28 and 2024-05-04, utilizing the array in the 6km (6A and 6D) and (H168) configuration. The observations were conducted at frequencies of GHz, GHz, GHz, GHz and GHz. The flux calibrator was the standard ATCA primary calibrator, 1934–638, and 1548+056 is used as phase calibrator. The observations are summarized in Table LABEL:table:atca.
The data reduction procedures were based on the Multichannel Image Reconstruction, Imaging Analysis and Display (miriad: Sault et al. 1995). The data reduction procedure was conducted in accordance with standard protocol. In order to mitigate the impact of radio frequency interference (RFI) during flux and phase calibration, the interactive flagging tool bflag was utilised. Furthermore, automated flagging routines pgflag were applied to address interference for the source. Any corrupted data identified during calibration was manually flagged.
Due to the limited ()-coverage and the low elevation of the source resulting in a very elongated beamshape, we were not able to perform self-calibration and cleaning. Fortunately, AT2019aalc was always dominating the flux within the resolution elements of our short observations. Therefore, we imaged the region around our source for each observation using an oversampling of the synthesised beam by a factor of at least 10. Under the assumption that AT2019aalc is always a point-source, we can take the highest pixel value in the image as our measured flux. Due to the uncertainties this method introduced we used a plain % flux error.
@@ 3.5 Spectroscopic observations
The Low Resolution Imaging Spectrograph (LRIS: Oke et al. 1995) mounted on the Keck-I telescope at Maunakea measured the spectrum of AT2019aalc, on 2021-07-06 (PI: Kulkarni). Later, during the second optical flare, we obtained more optical spectra of AT2019aalc using the following instruments: LRIS (PI: Kulkarni), DeVeny Spectrograph mounted on the Lowell Discovery Telescope (LDT) in Arizona (PI: Hammerstein) and the Double Beam Spectrograph (DBSP, Oke & Gunn 1982) mounted on the inch telescope at Palomar Observatory, California (PIs: Kulkarni and Kasliwal). The spectra were reduced following the standard procedures using the automated reduction pipeline of the LRIS (Perley 2019), the software package PypeIt (Prochaska et al. 2020) for DeVeny, and the software packages DBSP DRT (Mandigo-Stoba et al. 2022) and PypeIt (Prochaska et al. 2020) for the DBSP. Details of these observations of AT2019aalc are summarized in Table 2 while the spectra are shown in Fig. 17.
@@
To determine the type of AGN prior the flares, we performed a stellar population synthesis fit of the pre-flare SDSS spectrum with (pPXF: Cappellari & Emsellem 2004; Cappellari 2017, 2023) and the newest GALAXEV (Bruzual & Charlot 2003) models (CB2019), assuming a Kroupa (2001) initial mass function (IMF) and an upper mass limit of , as templates (see Fig. 18). Beforehand, the spectrum was corrected for Milky Way extinction using the dust reddening maps by Schlafly & Finkbeiner (2011) and Cardelli et al. (1989) extinction curve. Furthermore, it was redshift corrected and logarithmically rebinned. We included several emission lines to the fit and assumed four kinematic components: One for all stellar templates, one for all forbidden lines, one for all allowed emission lines, and a last one for a second, broad component of all Balmer lines up to . Additionally, we included two separate dust components, each one with the extinction curve of Cardelli et al. (1989): a first one for stellar, and a second one for nebular emission. We therefore fixed the Balmer line ratio and limited doublets of the preinstalled lines to their theoretical ratio using the keywords tie_balmer and limit_doublets of pPXF. However, we only fixed the ratio for the strongest lines of the Balmer series up to to avoid possible contamination of by the [Ne iii] line. For the fit with pPXF we included additive and multiplicative polynomials up to tenth order (degree=10, mdegree=10) to model the AGN continuum. In order to correct for intrinsic dispersion caused by the spectrograph, we assume a resolving power of at and at .555https://www.sdss4.org/dr17/spectro/spectro_basics/ Linear interpolation is used to calculate values for wavelengths in between.
@@
The newly observed spectra were fitted using the galaxy spectrum fitting tool PyQSOFit (Guo et al. 2018; Shen et al. 2019; Ren et al. 2024). The continuum of each spectrum was modeled with a combination of the following components: a power-law to represent the AGN continuum, polynomials to capture any residual smooth variations, and an Fe II template for the iron emission complex. Emission lines in the H and H complexes were modeled using a combination of broad and narrow Gaussian profiles to represent different kinematic components. The high-ionization coronal lines of interest were fitted with additional Gaussian profiles. The H complex shows a strong double-peaked feature around . This region is known as Bowen Fluorescence complex. It was fitted with a set of broad Gaussian components. We computed parameter uncertainties using MC resampling. Examples of the fitted spectra and the inferred emission line parameters are given in Fig. 19 and in Table 3, respectively.
@@ 4.1 Optical/UV
The optical and UV light curves of AT2019aalc are shown in the bottom panel of Fig. 3. The transient was first detected by the ZTF in January with an -band magnitude of mag. AT2019aalc brightened to a peak magnitude of () reached on 2019-06-18 on a timescale of days. The second flare started mid-May , approximately years after the initial flare (Veres et al. 2023). The second flare evolved very similarly to the original one. From the beginning of the second flare, a mag rise was seen up to the peak of the flare on 2023-07-13 with mag (), on a timescale of days. The second flare peaked at roughly times the peak luminosity of the first flare. The peak of the second flare was followed by a short plateau phase that lasted for around days. Both flares declined very slowly, on timescales of years. The first flare decayed with a power-law index of while the second with in the -band. The decaying phases are not monotonic, as bumps tend to appear. During the second optical flare, these bumps appear four times: end of December , mid-March , end of May and mid-August (see Figure 3). Both optical flares brightened with constant color () and showed reddening when decaying.
Prior to the discovery of AT2019aalc, the host galaxy of the transient was monitored by several optical surveys. To study the pre-discovery optical variability of AT2019aalc, we performed aperture photometry of the All-Sky Automated Survey for Supernovae (ASAS-SN: Shappee et al. 2014; Kochanek et al. 2017) and the Catalina Real-Time Transient Survey (CRTS: Djorgovski et al. 2011). In addition, Palomar Transient Factory (PTF Law et al. 2009) magnitudes were extracted from the PTF Lightcurve Table666https://irsa.ipac.caltech.edu. No significant optical variability is visible between April and the discovery of the transient in January by ZTF (see Fig. 4).
The UV monitoring of the source started on 2023-06-23 around days before the optical peak of the second flare. The U-band peak is coincident with the optical peak while the shorter wavelength bands peaked weeks later. We note that right after the -band peak, a short plateau phase started in optical and UV. The transient is most luminous at the shortest wavelength ranges covered by UVOT. The UV emission peaks around erg/s in the filter UW2. After the peak, the UW2U color shows reddening. Towards the end of the UVOT monitoring, the measured UM2 magnitudes decreased back to the level of the baseline GALEX-NUV magnitudes measured between and .
Comparing the peak UVM2 magnitude of the UVOT monitoring to the GALEX-DR5 (Bianchi et al. 2011) NUV magnitude implies that the UV luminosity increased by a factor of with respect to the archival flux. The amplitude of this flare is higher than typically seen for AGN. Only out of the AGN presented in the GALEX Time Domain Survey (Gezari et al. 2013) have larger amplitude NUV variability.

### v3

@@ 3.2 Infrared
The Palomar 200-inch (P200) near-infrared telescope equipped with the Wide Field Infrared Camera (WIRC, 2003SPIE.4841..451W) observed AT2019aalc on 2023-07-05. The imaging was obtained in the J-, H,- and Ks-bands. The data was processed with the Swarp software (2002ASPC..281..228B). We performed background subtraction using the photutils.background class and measured the flux using a circular aperture with a radius of pixels. The corresponding zero-point magnitudes for this aperture size were calculated using the automatic pipeline for each filter. The error bars are the computed Poisson and background error of the photometry222https://www.stsci.edu/hst/instrumentation/wfc3/data-analysis/photometric-calibration/ir-photometric-calibration and we also took into account the uncertainty of the derived zero-point magnitudes.
We obtained mid-infrared photometry from archival observations of the Wide-Field Survey Explorer (WISE, 2010AJ....140.1868W), which was continuously monitoring the sky at (W1) and, due to the NEOWISE project (2011ApJ...731...53M), at (W2). NEOWISE visits any point in the sky roughly every six months, with each visit consisting of about ten exposures. Using timewise (neckerJannisNeTimewiseV02024; 2025A&A...695A.228N), we downloaded all available multi-epoch photometry data from the AllWISE data release and single-epoch photometry from the NEOWISE-R data releases and stacked the data per visit to obtain robust measurements. All selected single-exposure photometry data points are within arcsecond of the nucleus of the host. The resulting light curve spans almost years, starting in February until July 2024.
@@ 3.3 X-ray
The X-ray Telescope (XRT, 2005SSRv..120..165B) observations, in the energy range of keV, were obtained simultaneously with the UVOT observations. The observations were performed in photon counting mode. We performed the data reduction using the Space Science Data Center (SSDC) interactive archive in the following way: first, cleaned event-files were produced using the xrtpipeline and xrtexpomap tasks. The ximage tool allowed us to determine the source detection level and its position. In case of non-detections, upper limits for the count rate were derived and converted to flux using WebPIMMS333https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl and assuming the mean power-law index of the observations that resulted in detections. When our target source was detected above the level, we ran swxrtdas in order to extract the source and the background spectra. Source counts were retrieved from a circle with a radius of pixel ( arcseconds), while background counts were extracted using an annular region with pixel inner/outer radius. Finally, the grppha tool of the heasoft 6.32 software package was used to perform background subtraction using the source and background spectra. We binned each spectrum to have a minimum of counts per spectral bin.
We fit the individual X-ray spectra using the XSPEC fitting environment (swift). First, we have determined the Galactic neutral hydrogen column density, , based on the HI4PI survey (2016A&A...594A.116H). The spectrum was then fitted using Cash-statistic (C-stat, 1976A&A....52..307C). To find the best-fitting model, we first fitted the stacked spectrum with a absorbed power-law (Tbabs*powerlaw in XSPEC) model, as the last XMM data of AT2019aalc before our monitoring program was best-fit with a power-law model (2023ATel16118....1P). When fitting the model, the galactic column density was fixed whilst the power-law index and the normalization were free to vary. The resultant fit quality is C-Stat for degrees of freedom. We then added a blackbody model component, to take into account a possible soft excess. The blackbody temperature and the normalization of the new component were also free to vary. We fit again, which resulted in a better fit with C-Stat for degrees of freedom. We implemented the likelihood ratio test (lrt in XSPEC) to estimate the statistical significance of the inclusion of the blackbody component. After running simulations, the test resulted only in smaller likelihood differences than the observed one, clearly implying that the difference is unlikely to be due to random fluctuations. Therefore, accounting for the soft X-ray excess results in a significantly better fit.
@@ 3.4.1 VLBI observations
We observed AT2019aalc with the European VLBI Network (EVN) and the enhanced Multi-Element Remotely Linked Interferometer Network (e-MERLIN, 2004SPIE.5489..332G). These observations were carried out at GHz (L-band) on 2023 June 09 (project code: EV027, PI: Veres). The observation time was hr. The experiment was performed in phase-referencing mode with -min duty cycles (with spending min on the target during each cycle) which resulted in an on-source time of h including fringe-finder scans and slewing. The data were recorded in left and right circular polarizations with a data rate of Gbit s-1. The correlation was done at the EVN Data Processor (JIVEcorrelator) at the Joint Institute for VLBI European Research Infrastructure Consortium (Dwingeloo, The Netherlands) with an averaging time of s. The Compact Symmetric Object (CSO) 15210430 (a.k.a. 4C 04.51) at an angular distance of was observed as phase-reference calibrator.
The data were calibrated with the U.S. National Radio Astronomy Observatory Astronomical Image Processing System (AIPS, aips) software package. The standard procedures including time flagging, amplitude calibration, correction for the dispersive ionospheric delays and global fringe-fitting of the calibrator sources were followed. We imaged the phase-calibrator with the Difmap software (difmap), using the hybrid mapping technique (hybridmapping). Using the clean component (CC) model we built up in Difmap, we repeated the fringe-fitting in AIPS for the phase-reference calibrator source to correct its structure. Finally, we interpolated the solutions of the last fringe-fitting to the data of our target source. We imaged the target source using the hybrid mapping technique in Difmap. At the end of the procedure, several steps of amplitude and phase self-calibration were performed. A simple elliptical Gaussian modelfit component (1995ASPC...82..267P) was fitted directly to the visibility data. Finally, to decrease the noise level, we performed clean iterations in the residual image with a small loop gain of .
@@ 3.4.2 ATCA monitoring
The Australia Telescope Compact Array (ATCA)444https://csiropedia.csiro.au/australia-telescope-compact-array/, located at Narrabri, New South Wales, Australia, was used to monitor AT2019aalc. The observations were conducted between 2023-06-28 and 2024-05-04, utilizing the array in the 6km (6A and 6D) and (H168) configuration. The observations were conducted at frequencies of GHz, GHz, GHz, GHz and GHz. The flux calibrator was the standard ATCA primary calibrator, 1934–638, and 1548+056 was used as phase calibrator. The observations are summarized in Table LABEL:table:atca.
The data reduction procedures were based on the Multichannel Image Reconstruction, Imaging Analysis and Display (miriad: miriad). The data reduction procedure was conducted in accordance with standard protocols. In order to mitigate the impact of radio frequency interference (RFI) during flux and phase calibration, the interactive flagging tool bflag was utilised. Furthermore, automated flagging routines pgflag were applied to address interference for the source. Any corrupted data identified during calibration was manually flagged.
Due to the limited ()-coverage and the low elevation of the source, which resulted in a very elongated beamshape, we were not able to perform self-calibration and cleaning. Fortunately, AT2019aalc always dominated the flux within the resolution elements of our short observations. Therefore, we imaged the region around our source for each observation using an oversampling of the synthesised beam by a factor of at least 10. Under the assumption that AT2019aalc is always a point-source, we can take the highest pixel value in the image as our measured flux. Due to the uncertainties this method introduced, we used a plain % flux error.
@@ 3.5 Spectroscopic observations
We observed the spectrum of AT2019aalc on 2021-07-06 using the Low Resolution Imaging Spectrograph (LRIS: 1995PASP..107..375O) mounted on the Keck-I telescope at Maunakea (PI: Kulkarni). Later, during the second optical flare, we obtained five more optical spectra of AT2019aalc using the following instruments: LRIS (PI: Kulkarni), DeVeny Spectrograph mounted on the Lowell Discovery Telescope (LDT) in Arizona (PI: Hammerstein) and the Double Beam Spectrograph (DBSP, 1982PASP...94..586O) mounted on the inch telescope at Palomar Observatory, California (PIs: Kulkarni and Kasliwal). The spectra were reduced following the standard procedures using the automated reduction pipeline of the LRIS (2019PASP..131h4503P), the software package PypeIt (2020JOSS....5.2308P) for DeVeny, and the software packages DBSP DRT (2022JOSS....7.3612M) and PypeIt (2020JOSS....5.2308P) for the DBSP. Details of these observations of AT2019aalc are summarized in Table 2 while the spectra are shown in Fig. 17.
@@
To determine the type of AGN prior to the flares, we performed a stellar population synthesis fit of the pre-flare SDSS spectrum with (pPXF: 2004PASP..116..138C; 2017MNRAS.466..798C; 2023MNRAS.526.3273C) and the newest GALAXEV (2003MNRAS.344.1000B) models (CB2019), assuming a 2001MNRAS.322..231K initial mass function (IMF) and an upper mass limit of , as templates (see Fig. 18). Beforehand, the spectrum was corrected for Milky Way extinction using the dust reddening maps by 2011ApJ...737..103S and the 1989ApJ...345..245C extinction curve. Furthermore, it was redshift corrected and logarithmically rebinned. We included several emission lines in the fit and assumed four kinematic components: one for all stellar templates, one for all forbidden lines, one for all allowed emission lines, and a last one for a second, broad component of all Balmer lines up to . Additionally, we included two separate dust components, each one with the extinction curve of 1989ApJ...345..245C: a first one for stellar, and a second one for nebular emission. We therefore fixed the Balmer line ratio and limited doublets of the preinstalled lines to their theoretical ratio using the keywords tie_balmer and limit_doublets in pPXF. However, we only fixed the ratio for the strongest lines of the Balmer series up to to avoid possible contamination of by the [Ne iii] line. For the fit with pPXF we included additive and multiplicative polynomials up to tenth order (degree=10, mdegree=10) to model the AGN continuum. In order to correct for intrinsic dispersion caused by the spectrograph, we assume a resolving power of at and at .555https://www.sdss4.org/dr17/spectro/spectro_basics/ Linear interpolation is used to calculate values for wavelengths in between.
@@
The newly observed spectra were fitted using the galaxy spectrum fitting tool PyQSOFit (qso1; qso2; qso3). The continuum of each spectrum was modeled with a combination of the following components: a power-law to represent the AGN continuum, polynomials to capture any residual smooth variations, and an Fe II template for the iron emission complex. Emission lines in the H and H complexes were modeled using a combination of broad and narrow Gaussian profiles to represent different kinematic components. The high-ionization coronal lines of interest were fitted with additional Gaussian profiles. The H complex shows a strong double-peaked feature around . This region is known as Bowen Fluorescence complex. It was fitted with a set of broad Gaussian components. We computed parameter uncertainties using MC resampling. Examples of the fitted spectra and the inferred emission line parameters are given in Fig. 19 and in Table 3, respectively.
@@ 4.1 Optical/UV
The optical and UV light curves of AT2019aalc are shown in the bottom panel of Fig. 3. The transient was first detected by the ZTF in January with an -band magnitude of mag. AT2019aalc brightened to a peak magnitude of () reached on 2019-06-18 on a timescale of days. The second flare started mid-May , approximately four years after the initial flare (re-br). The second flare evolved very similarly to the original one. From the beginning of the second flare, a mag rise was seen up to the peak of the flare on 2023-07-13 with mag (), on a timescale of days. The second flare peaked at roughly times the peak luminosity of the first flare. The peak of the second flare was followed by a short plateau phase that lasted for around days. Both flares declined very slowly, on timescales of years. The first flare decayed with a power-law index of while the second with in the -band. The decaying phases are not monotonic, as bumps tend to appear. During the second optical flare, these bumps appear four times: end of December , mid-March , end of May and mid-August (see Figure 3). Both main optical flares brightened with constant color () and showed reddening when decaying.
Prior to the discovery of AT2019aalc, the host galaxy of the transient was monitored by several optical surveys. To study the pre-discovery optical variability of AT2019aalc, we performed aperture photometry of the All-Sky Automated Survey for Supernovae (ASAS-SN: 2014ApJ...788...48S; 2017PASP..129j4502K) and the Catalina Real-Time Transient Survey (CRTS: 2011arXiv1102.5004D). In addition, Palomar Transient Factory (PTF 2009PASP..121.1395L) magnitudes were extracted from the PTF Lightcurve Table666https://irsa.ipac.caltech.edu. No obvious signs of flaring activity is visible between April and the discovery of the transient in January by ZTF (see Fig. 4).
The UV monitoring of the source started on 2023-06-23 around days before the optical peak of the second flare. The U-band peak is coincident with the optical peak, while the shorter wavelength bands peaked weeks later. Furthermore, the U- and UVW2-band peaks occur closer in time than the UVW1 and UVM2 peaks, suggesting a complex SED evolution near the maximum. We note that right after the -band peak, a short plateau phase started in optical and UV. The transient is most luminous at the shortest wavelength ranges covered by UVOT. The UV emission peaks around erg/s in the filter UVW1. After the peak, the UVW2U color shows reddening. Towards the end of the UVOT monitoring, the measured UVM2 magnitudes decreased back to the level of the baseline GALEX-NUV magnitudes measured between and .
Comparing the peak UVM2 magnitude of the UVOT monitoring to the GALEX-DR5 (2011Ap&SS.335..161B) NUV magnitude implies that the UV luminosity increased by a factor of with respect to the archival flux. The amplitude of this flare is higher than typically seen for AGN. Only out of the AGN presented in the GALEX Time Domain Survey (Gezari13) have larger amplitude NUV variability.